Ship reports banner with links Home IODP-USIO daily/weekly reports

< July | September >

IODP Expedition 327

Juan de Fuca Ridge-Flank Hydrogeology

Daily Science Report for 31 August 2010

LOCATION: Hole U1363A (Proposed Site GRB-1A; 47°17.3555'N, 128°2.1107'W; water depth 2689 m)
Hole U1363B (Proposed Site GRB-1A; 47°17.3518'N, 128°2.1060'W; water depth 2690 m)

SCIENCE UPDATE: We continued drilling without coring in Hole U1363A and reached basement at 2747 m (58 m below seafloor) by 0130 hr. The extended core barrel (XCB) bit was recovered at 0300 hr and the drill string was pulled out of the seafloor at 0315 hr, ending operations at Hole U1363A. Hole U1363B was spudded at 0530 hr. Advanced piston coring (APC) Cores 1H through 6H advanced to 42.5 mbsf and were recovered between 0550 and 1300 hr. Four APCT3 temperature measurements were taken with Cores 3H-6H, but the fourth attempt was unsuccessful due to limited penetration into turbidites. Following APC refusal at 42.5 mbsf, we switched to XCB coring. Cores 7X and 8X advanced to 55 mbsf and were recovered at 1450 and 1610 hr. The SET temperature tool was deployed following Core 8X and was on deck by 1800 hr. Core 9X advanced across the sediment/basement interface to 57 mbsf. Core 10X advanced another ~4 m and was on deck at 2100 hr. The drill string was pulled out of the seafloor at 2140 hr, ending operations at Hole U1363A. The ship offset in dynamic positioning mode to Hole U1363C. Hole U1363C was spudded at 2255 hr and drilling without coring had advanced to 2728 m (39 mbsf) by midnight.
Core description, shipboard sampling, and lab measurements started on the Site U1363 cores. Review of Site U1362 reports continued. Weather on site: wind 8 knots gusting to 10 knots, seas 4 ft, swells 9 ft, air temperature 15.5°C, max. heave 1.2 m, max. pitch 1.5°, max. roll 0.75°.


Daily Science Report for 30 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)
Hole U1363A (Proposed Site GRB-1A; 47°17.354'N, 128°2.112'W; water depth 2689 m)

SCIENCE UPDATE: The VIT was run to the seafloor with a grappling hook and the beacon was grabbed at 0146 hr. The beacon was on deck by 0255 hr and it was discovered that the beacon had not released properly because it had a tangled tether, which kept the weight attached. The drill string was pulled up by 0815 hr and the rig was secured for transit by 1300 hr. The thrusters were raised at 1315 hr and the ship was underway for Grizzly Bare Site U1363 (Scientific Prospectus Site GRB-1A). The thrusters were lowered at 1623 hr and the positioning beacon was deployed at 1700 hr. An Advanced Piston Coring (APC)/Extended Core Barrel (XCB) bottom-hole assembly was made up by 1830 hr and the drill string was run to the seafloor. Seafloor was tagged at 2689 m at 2250 hr. Hole U1363A was spudded at 2300 hr and by midnight we had drilled down to 2725 m (36 m below seafloor).
Review of Site U1362 reports continued. A measurement plan was developed for the Grizzly Bare sites. Weather on site: wind 12 knots gusting to 15 knots, seas 3 ft, swells 6 ft, air temperature 16°C, max. heave 1 m, max. pitch 1.5°, max. roll 1.5°.


Daily Science Report for 29 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: The drill string with the CORK assembly was run to 2658 m by 0200 hr. The instrument string with osmosamplers and temperature loggers was deployed between 0200 and 0600 hr. The instrument string landed in the CORK at 0630 hr. Hole U1362B was reentered with the CORK assembly for the eleventh time at 0648 hr. Shortly after water was observed flowing onto the rig floor timed with the ship’s heave, raising concerns that perhaps the instrument string was not properly seated. The drill string was pulled out of the reentry cone and the instrument string was pulled out of the hole at 0945 hr. The instrument string latch and Spectra rope were inspected and found to be OK. The instrument string was redeployed and landed in the CORK head by 1315 hr. Hole U1362B was reentered at 1330 hr and the CORK assembly landed in the reentry cone at 1515 hr. The inflatable packers were pressured up at 1000 psi for 30 min. The CORK platform was attached with slings to the Vibration Isolated Television (VIT) frame and deployed through the moonpool at 1800 hr. The platform was released at 1930 hr and the VIT was recovered so the slings could be removed. The VIT was redeployed at 2100 hr and the CORK running tool was released at 2217 hr, completing the Hole U1362B CORK installation.
Review of Site U1362 reports continued. CORK specialists assisted with the instrument string assembly. A meeting was held to discuss end of expedition and postcruise activities. The outreach team gave an overview of all the projects they have been working on. Coring near the Grizzly Bare outcrop is expected to start tomorrow. Weather on site: wind 5 knots gusting to 8 knots, seas 3 ft, swells 5 ft, air temperature 14°C, max. heave 1.2 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 28 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: The drill string was pulled out of the hole and cleared the rotary table at 0510 hr. The Hole U1362B L-CORK was assembled between 0530 and 2245 hr and comprised a bull nose, 3 perforated collars, a joint of perforated 5-1/2 inch casing, an inflatable 4-1/2 inch packer, a set of swellable 4-1/2 inch packers, 19 joints of 4-1/2 inch casing, and the CORK head. The drill string with the CORK assembly was run to 1175 m by midnight.
Review of Site U1362 reports continued. CORK specialists assisted with the CORK assembly. Weather on site: wind 22 knots gusting to 28 knots, seas 10 ft, swells 5 ft, air temperature 17°C, max. heave 2 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 27 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: The tracer injection experiment in Hole U1362B continued with the following stages: 1) sea water was pumped at 20 strokes per minute (spm) from 0000 to 0500 hr using the rig floor pump and the second salt tracer (Cs and Ho) was injected at 30 spm at 0400 hr using the cement pump, 2) fresh water was pumped at 20 spm from 0500 to 0600 hr and microsphere tracers were injected at 30 spm for 9 minutes using the cement pump, 3) sea water was pumped at 20 spm from 0600 to 0900 hr and bacteria tracers were injected at 30 spm for 9 minutes using the cement pump, 4) pumping ceased from 0900 to 1000 hr to allow the hole to equilibrate. The running tool was pulled from the casing hanger, the reentry cone was observed for signs of back flow with none seen, and the drill string was pulled out of the hole at 1100 hr. The osmosamplers were recovered at 1600 hr and a drilling bottom-hole assembly was made up for a depth check and potential conditioning of the hole. The drill string was run to bottom and Hole U1362B was reentered for the tenth time at 2130 hr. The drill string was advanced without rotation or circulation until it tagged bottom at 3020 m (348 m below seafloor).
Review of Site U1362 reports continued. A meeting was held to discuss the sampling plan for the Grizzly Bare sites. Outreach officer Stephanie Keske gave a presentation on computer graphics animation and showed the CORK project she has been working on. Weather on site: wind 10 knots gusting to 12 knots, seas 3 ft, swells 5 ft, air temperature 16°C, max. heave 0.6 m, max. pitch 0.5°, max. roll 0.5°. Weather is expected to deteriorate over the next two days.


Daily Science Report for 26 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: The drill string with the injection assembly was lowered to 2644 m by 0300 hr. The subsea TV signal was lost at 0300 hr and the TV was brought back to the surface and repaired by 0330 hr. Circulation rates were recorded in preparation for the 24-hr injection test. Hole U1362B was reentered for the ninth time at 0642 hr. The drill string with the injection assembly was lowered to 2926 m (255 m below seafloor) and the running tool landed in the casing hanger at 0800 hr. The tracer injection experiment started at 0800 hr and consisted of the following stages: 1) the first hour was used to establish baseline conditions without pumping, 2) sea water was pumped at 20 strokes per minute (spm) from 0900 to 1000 hr using the rig floor pump, 3) fresh water was pumped at 20 spm from 1000 to 1100 hr, 4) sea water was pumped at 20 spm from 1100 to 1200 hr, 5) sea water was pumped at 20 spm from 1200 to 1300 hr and the first salt tracer (Cs and Er) was injected at 30 spm from 1200 to 1209 hr using the cementing pump, 6) sea water was pumped at 20 spm from 1300 to 2400 hr. The experiment is scheduled to end at 1000 hr tomorrow. Throughout the experiment scientists working in shifts collected samples of the injected fluid at 15 min, 30 min, or 1 hr depending on sample type.
Review of Site U1362 reports continued. Outreach officers and scientists participated in a videoconference with the Harte Research Institute for Gulf of Mexico Studies at Texas A&M, Corpus Christi. Weather on site: wind 12 knots gusting to 15 knots, seas 4 ft, swells 5 ft, air temperature 16°C, max. heave 0.6 m, max. pitch 0.5°, max. roll 0.5°.


Daily Science Report for 25 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: Between 0000 and 1115 hr, Hole U1362B was swept with mud, washed, and reamed from 2944 to 3011 m (279-339 m below seafloor) four consecutive times before waiting for 1 hr for hole to stabilize. The accumulated fill was washed and reamed to total depth and the hole was swept with mud three more times by 1530 hr. The drill string was pulled out of the hole and cleared the rotary table at 2100 hr. The osmosamplers were loaded, the tracer injection bottom-hole assembly was completed by 2200 hr, and the drill string had reached 794 m by midnight.
Scientists finalized measurements and worked on their Site U1362 reports. Weather on site: wind 16 knots gusting to 18 knots, seas 3 ft, swells 4 ft, air temperature 16°C, max. heave 0.6 m, max. pitch 0.5°, max. roll 0.5°.


Daily Science Report for 24 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: Hole U1362B was reentered for the eighth time at 0125 hr. The hole was repeatedly washed and reamed to 3011 m (339 m below seafloor) and swept with mud. Between 1315 and 2400 hr, the 9-7/8 inch hole was extended to 3031 m (359 mbsf).
Lab measurements continue. Scientists turned in their Site U1362 reports. The imaging specialist recorded end-of-cruise interviews with some of the outreach officers. Designs were submitted for the t-shirt logo. Weather on site: wind 6 knots gusting to 9 knots, seas 3 ft, swells calm, air temperature 16°C, max. heave 0.5 m, max. pitch 0.25°, max. roll 0.25°.


Daily Science Report for 23 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: The drill string with the tracer injection test bottom-hole assembly (BHA) was lowered to 2086 m by 0200 hr. A pre-injection dry run of the circulation, tracer injection, and sampling procedures was conducted between 0330 and 0630 hr. Hole U1362B was reentered for the seventh time at 0715 hr. The drill string was lowered to 2935 m (263 m below seafloor) and had to be worked down until the running tool landed in the casing hanger. The 24-hr tracer injection test started at 0900 hr by recording baseline conditions. At 1000 hr, the injection test began with 20 strokes per minute on the rig mud pump. It was quickly discovered that the pump pressure was too high and the injection sub was pulled back inside the casing by 1030 hr. Circulation with normal pressures was reestablished, the running tool was landed again, but pump pressure again became too high by 1130 hr. The consensus was that there was a problem with the hole and the decision was made to pull out. The drill string was pulled out and the BHA cleared the rotary table at 1700 hr. A drilling BHA with a 9-7/8 inch drill bit was made up and the drill string was run to 2644 m by 2330 hr.
Core description and lab measurements continue. Scientists have been working on the Site U1362 reports. The imaging specialist recorded end-of-cruise interviews with some of the outreach officers. Weather on site: wind light, seas calm, swells calm, air temperature 17°C, max. heave 0.7 m, max. pitch 0.25°, max. roll 0.25°.


Daily Science Report for 22 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: Hole U1362B was drilled to a total depth of 3011 m (339 m below seafloor) by 0930 hr. The hole was then washed and reamed and swept with mud between 0930 and 1200 hr. The first wiper trip between 2944 and 3011 m ended at 1415 hr. The second wiper trip ended at 1600 hr. The hole was swept with mud again and the drill string was pulled out of the hole between 1700 and 2200 hr. At 2200 hr, the crew started making up the injection test bottom-hole assembly.
Core description and lab measurements continue. An operations meeting was held to discuss the tracer injection experiment. A videoconference was held with the R/V Thompson as it was returning from Juan de Fuca Ridge. Weather on site: winds 6 knots gusting to 10 knots, seas calm, swells calm, air temperature 15°C, max. heave 1.4 m, max. pitch 0.25°, max. roll 0.25°.


Daily Science Report for 21 August 2010

LOCATION: Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: Hole U1362B was reentered at 0423 hr. The drill string was run to 2844 m (172 m below seafloor) before encountering fill. The hole was reamed to 2934 m (262 mbsf), the cement plug was drilled out, and the 14-3/4 inch hole was deepened to the original total depth of 2954 m (282 mbsf). The hole was swept with mud and a 9-7/8 inch hole was drilled to 3005 m (333 mbsf) by midnight.
Core description and lab measurements continue. Scientists, outreach officers, technical staff, and crew members participated in the Post-2013 Drilling Program Naming Workshop. A science meeting was held to discuss operations at Hole U1362B and possibly the alternate Grizzly Bare sites. CORK specialists, engineers, and technical staff formulated a plan for the upcoming 24-hr tracer injection experiment. Weather on site: winds 8 knots gusting to 12 knots, seas 3 feet, swells 4 feet, air temperature 15°C, max. heave 0.4 m, max. pitch 0.5°, max. roll 0.5°.


Daily Science Report for 20 August 2010

LOCATION: Hole 1027C (Leg 168; 47° 45.387'N, 127° 43.867; water depth 2667 m)
Hole U1362B (Proposed Site SR-2; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: Work on fabricating a new recovery tool for the Hole 1027C CORK continued until 0600 hr. The recovery tool was deployed at 0630 hr, the CORK head was engaged at 1200 hr, and the lower set of latches had been engaged by 1245 hr. Between 1245 and 1600 hr,several attempts were made to release the CORK, eventually pulling up to 100,000 lbs. At 1600 hr it was decided to terminate efforts to recover the CORK as the latching mechanism appeared to be stuck. The CORK head was released at 1615 hr, the beacon was recovered at 1741 hr, and the ship started moving in dynamic positioning mode to Hole U1362B. The drill string cleared the rotary table at 2130 hr, ending operations at Hole 1027C. A tricone bit drilling bottom-hole assembly was made up by 2300 hr, the ship arrived at Hole U1362B at 2312 hr, and the drill string was run to 738 m by midnight.
Core description and lab measurements continue. A videoconference was held with a University of Southern California Wrigley Institute science camp. Weather on site: winds 9 knots gusting to 12 knots, seas 4 feet, swells 4 feet, air temperature 14.5°C, max.heave 0.5 m, max. pitch 0.5°, max. roll 0.5°.


Daily Science Report for 19 August 2010

LOCATION: Hole 1027C (Leg 168; 47° 45.387'N, 127° 43.867; water depth 2667 m)

SCIENCE UPDATE: Work continued on fabricating a new recovery tool for the Hole 1027C CORK.
Core description and lab measurements continue. Outreach Officer Jackie Kane led a videoconference with her students at St. Ursula's Academy in Toledo, Ohio. The ROV challenge was held in the morning and at night. Weather on site: winds 12 knots gusting to 15 knots, seas 4 feet, swells 5 feet, air temperature 14.5°C, max. heave 0.6 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 18 August 2010

LOCATION: Hole U1301B (Expedition 301; 47° 45.228'N, 127° 45.827'W; water depth 2667 m)
Hole 1027C (Leg 168; 47° 45.387'N, 127° 43.867; water depth 2667 m)

SCIENCE UPDATE: The temperature profile at Hole U1301B continued until bottom was tagged at 3037 m (370 m below seafloor) and the temperature loggers were pulled out of the hole at 0145 hr. The Electronic Release System (ERS) was rigged up to the Schlumberger electric line, the replacement instrument string was deployed, and it was released with the ERS at 0530 hr. The drill string was pulled out of the hole and the ship was offset in dynamic positioning mode to Hole 1027C. The ship was maneuvered and the CORK head was engaged with the recovery tool at 1015 hr. Between 1015 and 1345 hr several attempts were made to pull the CORK but it appeared to be stuck with the latches holding it in place not releasing. Following discussions of possible options, it was determined that the recovery tool was not capable of reaching and engaging the lower set of latches. At 1830 hr, the drill string was pulled out of the hole, the recovery tool was retrieved, and work began on fabricating a modified recovery tool.
Core description and lab measurements continue. A science meeting was held to discuss operations at Hole 1027C. Outreach Officer Jackie Kane led a videoconference with the Toledo (Ohio) Imagination Station. The outreach team and several scientists participated in a costume making workshop led by Katie Inderbitzen. Weather on site: winds 21 knots gusting to 25 knots, seas 5 feet, swells 4 feet, air temperature 14°C, max. heave 0.6 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 17 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)
Hole U1301B (Expedition 301; 47° 45.228'N, 127° 45.827'W; water depth 2667 m)

SCIENCE UPDATE: The subsea TV and suspended CORK platform reached bottom at 0230 hr and the platform was released at 0330 hr after the “dogs” holding it to the VIT frame did not release simultaneously. The subsea TV was pulled out of the hole to remove the platform slings and when it was run back to bottom, it was discovered that the platform was resting unevenly on the CORK fins. The platform was nudged with the VIT frame and fell into its proper position on top of the reentry cone at 0515 hr. The running tool was released from the CORK head at 0600 hr, completing the successful installation of the Hole U1362A L-CORK. The beacon was released at 0720 hr, and was recovered at 0816 hr. While the ship was offset to Hole U1301B, the beacon had its batteries changed and was redeployed midway between Holes U1362A and U1301B at 1007 hr. The CORK running tool cleared the rig floor at 1045 hr, ending operations at Hole U1362A and starting operations at Hole U1301B. The CORK recovery tool was made up, the subsea TV was deployed, and the drill string was run to bottom. The ship maneuvered over Hole U1301B and the CORK recovery tool engaged with the CORK head at 1730 hr. After repeated attempts to free the instrument string the line parted at 1920 hr. The drill string was pulled out of the hole and the CORK recovery tool was disengaged. The recovered partial instrument string consisted of ~377 m of Spectra rope and five temperature loggers. The drill string was run back into the hole for a depth check of the remaining hole and to obtain a temperature profile. The CORK head was reengaged at 2225 hr and the temperature loggers were run into the hole.
Core description and lab measurements continue. Weather on site: winds 16 knots gusting to 18 knots, seas 6 feet, swells 4 feet, air temperature 14°C, max. heave 0.8 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 16 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: By 0200 hr, the CORK had been pulled up to the moonpool. The running tool was removed, the instrument string was unlatched, the Spectra rope was shortened by 10.2 m and the instrument string was redeployed in the drill pipe at 0600 hr. The drill string was slowly rigged into the hole, filling every stand with water, and was close to bottom by 1115 hr. Hole U1362A was reentered for the eleventh time at 1150 hr, the drill string with the CORK assembly was deployed to 3142 m (470 m below seafloor), and the L-CORK landed at 1450 hr. The packers were set and pressured up and the subsea TV was recovered by 1600 hr. The CORK platform was suspended from the subsea TV frame and deployed through the moonpool at 1800 hr. At 2007 hr and 2540 m depth, the subsea camera failed and had to be recovered. Following repairs, the subsea TV and CORK platform were redeployed at 2345 hr.
Core description and lab measurements continue. The Expedition 327 international kite competition was held in the afternoon. Weather on site: winds 14 knots gusting to 18 knots, seas 4 feet, swells 3 feet, air temperature 14°C, max. heave 0.2 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 15 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: The remaining chemistry and pressure miniscreens were installed by 0145 hr, followed by 22 joints of 4-1/2 inch casing with the umbilicals secured on each joint. The L-CORK head was attached to the upper joint of casing at 0945 hr. The final umbilical terminations were made and attached to the casing by 1115 hr. The CORK valves were opened to release trapped air, the CORK was lowered into the water and pulled back up, and the valves were closed. Between 1300 and 1630 hr, the drill string was run a little short of the seafloor. Between 1630 and 1815 hr, the osmotic samplers and sinker bar weights were deployed into the drill pipe, followed by the thermistors on the Spectra rope. The instrument string was slowly lowered into the CORK, landed, and latched by 1945 hr. Before reentry, an inspection with the underwater camera showed that the instrument string extended below the end of the CORK, preventing us from deploying the CORK. At 2230 hr, the decision was made to pull the CORK back up.
Core description and lab measurements continue. The Curator worked with the science party to finalize the location of personal samples before taking them. The weekly fire and boat drill was followed by group photos. Weather on site: winds 6 knots gusting to 8 knots, seas 2 feet, swells 3 feet, air temperature 18°C, max. heave 0.2 m, max. pitch 0.5°, max. roll 0.5°.


Daily Science Report for 14 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: At 0230 hr, the decision was made to modify the bottom swellable packer. For the next several hours, the outer diameter of the packer was turned down from 9-3/4 to 8-1/2 inches. At the same time, the microbiology, chemistry, and pressure umbilicals were connected to the top of the lower (open hole) inflatable packer and the microbiology, chemistry, and pressure miniscreens were installed below the inflatable packer. The first set of modified swellable packer halves was installed at 1230 hr and the umbilical lines running between the inflatable and swellable packers were secured with metal bands. The modification of the second set of packer halves took until 1830 hr to complete and they were installed by 1930 hr. Eight joints of 4-1/2 inch casing were added to the drill string by 2200 hr. By midnight, the second pair of inflatable and swellable packers had been added, all umbilical connections had been terminated, and the remaining miniscreens had been installed.
Core description and lab measurements continue. The Sample Allocation Committee reviewed all samples requested from the working half sections. Weather on site: winds 10 knots gusting to 13 knots, seas 2 feet, swells 4 feet, air temperature 18°C, max. heave 0.4 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 13 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: The Hole U1362A packer flow test at 3096.5 m (424.5 m below seafloor) was completed by 0430 hr with good data recovered. The drill string was raised to 2989 m (317 mbsf) for the second packer test inside the 10-3/4 inch casing. The packer was set but could not lock-in pressure. After an inspection of the setting go-devil found no obvious problem, the decision was made at 1000 hr to suspend the packer test. The drill string was pulled out of the seafloor, the packer assembly was laid out, and the recovered packer element was found to be severely gouged and leaking water. The bit cleared the rig floor at 1800 hr and the crew prepared to run the 4-1/2 inch casing for the bottom of the CORK assembly. Three 8-1/4 inch perforated/coated drill collars were made up and run in the hole by 1945 hr. These were followed by a joint of 5-1/2 inch perforated/coated casing, the lower inflatable packer, a 4-1/2 inch casing collar with the umbilical landing seat, and the lower swellable packer intended to be placed in the open hole. Concerns were raised about the diameter of the swellable packer (9-3/4 inch for the rubber element) and mitigation discussions were initiated.
Core description and lab measurements continue. A sampling party was held to identify samples for postcruise analyses. Weather on site: winds 10 knots gusting to 13 knots, seas 6 feet, swells 5 feet, air temperature 16°C, max. heave 0.8 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 12 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Hole U1362A was reentered for the tenth time at 0211 hr. The Schlumberger logging tool suite was made up by 0545 hr. The drill string was lowered to 3178 m (506 m below seafloor) and the first logging pass to 2980 m (308 mbsf) was completed by 0945 hr. The drill string was lowered to 3178 m and the second logging pass to 2980 m (308 mbsf) was completed by 1115 hr. The drill string was lowered to 3178 m and the third logging pass to 3145 m (473 mbsf) was completed by 1230 hr. The interval for packer flow testing was identified at 3090-3102 m (418-430 mbsf) based on the caliper (hole diameter) log. The logging tools were pulled out of the hole and rigged down by 1615 hr. A depth check was conducted with the logging/packer bottom-hole assembly and the bottom of the hole was tagged at 519 mbsf. With the CORK assembly planned to extend to 470 mbsf, there will be a 49 m "rat" hole below the CORK to catch debris. The drill string was raised to 3102 m (430 mbsf) in preparation for the packer flow test. The center of the packer was positioned at 3096.5 m (424.5 mbsf) at 2000 hr. Initially problems were encountered trying to set the packer with ship's heave being the most likely cause. Eventually the packer set and locked-in pressure in the desired depth interval, allowing the packer flow test to begin.
Core description and lab measurements continue. Held science meeting to discuss preliminary data from Site U1362 and coordinate sampling party. The two French teachers and several scientists participated in a videoconference with students in Nouméa (New Caledonia). Weather on site: winds 26 knots gusting to 30 knots, seas 12 feet, swells 4 feet, air temperature 15°C, max. heave 1.6 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 11 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: A third wiper trip was conducted in Hole U1362A between midnight and 0400 hr and the hole was swept with mud. The drill string was raised to 308 m below seafloor (mbsf), inside the casing shoe, and kept on standby from 0600 to 0930 hr to let the hole stabilize. Between 0930 and 1045 hr, the drill pipe was lowered without rotation or circulation for a depth check and tagged fill at 512 mbsf, with the bottom of the CORK string planned for 476 mbsf. The hole washed and reamed for the last time to 528 mbsf and swept with mud. The drill string was raised and cleared the rotary table at 1845 hr. In preparation for logging, the wireline logging/packer bottom-hole assembly was made up by 1945 hr. The drill pipe was run into the hole and had reached 2034 m by midnight.
Core description, shipboard sampling, and lab measurements continue. Outreach activities continued with a discussion on web site metrics. Weather on site: winds 31 knots gusting to 33 knots, seas 10 feet, swells 8 feet, air temperature 14°C, max. heave 1.5 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 10 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Hole U1362A was drilled ahead with a 9 7/8 inch tri-cone bit to 528 m below seafloor by 1700 hr, with the rate of penetration averaging to 2 m/hr. The hole was repeatedly swept with mud and had two wiper trips conducted from the 10-3/4 inch casing shoe to total depth.
Core description, shipboard sampling, and lab measurements continue. Outreach activities continued with a microbe culturing activity and a talk on seafloor and ocean crust bacteria. Weather on site: winds 20 knots gusting to 22 knots, seas 6 feet, swells 5 feet, air temperature 14.5°C, max. heave 1.5 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 9 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: In preparation for logging operations and the CORK deployment, Hole U1362A was conditioned with a second wiper trip and mud sweep by 0230 hr. A third wiper trip and mud sweep was completed by 0730 hr. The amount of fill encountered at the bottom of the hole each time led to the decision to deepen the hole. The drill string was pulled out of the seafloor and cleared the rig floor at 1400 hr. A drilling assembly with a 9 7/8 inch tri-cone bit was made up, the drill string was lowered to the seafloor, and Hole U1362A was reentered for the ninth time at 1935 hr. The bit reached 476 m below seafloor (mbsf) at 2245 hr and had washed and reamed to 496 mbsf by 2330 hr.
Core description, shipboard sampling, and lab measurements continue. Engineers, CORK specialists, and technical staff held a meeting to discuss CORK operations and safety issues. Outreach officers, scientists, and technical staff participated in a videoconference with a science camp at University of Southern California. The ROVs were deployed in the pool at night to test their lights. Weather on site: winds 15 knots gusting to 18 knots, seas 4 feet, swells 5 feet, air temperature 14°C, max. heave 0.5 m, max. pitch 1°, max. roll 1°. The weather is expected to deteriorate in the next several days.


Daily Science Report for 8 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Cores U1362A-19R, 20R, and 21R arrived on deck at 0350, 0915, and 1255 hr, respectively, advancing to 3168 m (496 m below seafloor). Recovery averaged 26%. To prepare for logging and CORK operations, the hole was conditioned with two wiper trips starting at 1300 hr.
Core description, shipboard sampling, and lab measurements continue. Logging staff are preparing for logging operations. Engineering, science, and technical staff are preparing for CORK operations. Weather on site: winds 20 knots gusting to 24 knots, seas 6 feet, swells nil, air temperature 14°C, max. heave 0.8 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 7 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Cores U1362A-16R and 17R arrived on deck at 0145 and 0920 hr, respectively, advancing to 463 m below seafloor (mbsf). The slow rotation rate continued and the combined recovery for these cores was 8.95 m or 93.2%. Core 18R advanced to 472 mbsf and was recovered at 1745 hr. Although recovery remained high at 87.2%, the rotation rate increased at the bottom of the interval, indicating a return to a less massive formation. The pipe got stuck at 1745 hr and was freed by 2045 hr. Between 2045 and 2300 hr, the hole was washed and reamed to 472 mbsf.
Core description, shipboard sampling, and lab measurements continue. The outreach officers with the assistance of Geoff Wheat and Sam Hulme, deployed the ROVs they constructed in the pool. Weather on site: winds 18 knots gusting to 22 knots, seas 6 feet, swells 5 ft, air temperature 14°C, max. heave 0.4 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 6 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Drill pipe continued to be deployed and the bottom of Hole U1362A was tagged at 393 m below seafloor (mbsf) at 0300 hr. Between 0300 and 1230 hr, the hole was washed and reamed to 443 mbsf. Core 15R arrived on deck at 1745 hr with 21.5% recovery. By midnight, the cutting of Core 16R had advanced to 458 mbsf. Formation appears to be more massive at this depth.
Core description, shipboard sampling, and lab measurements continue. Scientists and crew painted the CORK bodies under the leadership of outreach officer Dinah Bowman. Weather on site: winds 18 knots gusting to 22 knots, seas 6 feet, swells 5 ft, air temperature 15°C, max. heave 0.3 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 5 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Core U1362A-14R advanced to 448.6 m below seafloor (mbsf) and was recovered at 0530 hr. Given the time the bit has been used and the additional coring expected at this hole, the decision was made to change bits. The drill string was pulled out of the hole and the bit cleared the rotary table at 1200 hr. The coring assembly was inspected and significant wear was discovered on the upper side of the bit as a result of the repeated reaming of the hole. The bit was changed and maintenance was conducted on the heave compensator laser tracking device to improve the RigWatch depth tracking capability. The drill string was lowered to 2022 m, the ship was maneuvered, and Hole U1362A was reentered for the eighth time at 2335 hr.
Core description, shipboard sampling, and lab measurements continue. The two French teachers and several scientists participated in a videoconference with students in Nouméa (New Caledonia). Weather on site: winds 20 knots gusting to 24 knots, seas 6 feet, swells 7 ft, air temperature 13°C, max. heave 0.9 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 4 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Core U1362A-10R advanced to 415 m below seafloor (mbsf) and was recovered at 0330 hr. The hole was washed, reamed, and swept with mud and coring resumed at 0630 hr. Core 11R advanced to 420 mbsf and was recovered at 0920 hr. The pipe got stuck while trying to recover Core 12R, which arrived on deck at 1600 hr. The hole was again washed, reamed, and swept with heavy mud to improve conditions. The next core barrel was deployed at 1845 hr with a target depth of 439 mbsf.
Core description, shipboard sampling, and lab measurements continue. Several scientists and outreach officers participated in a videoconference with the Smithsonian Institution (Washington, DC). Weather on site: winds 18 knots gusting to 22 knots, seas 6 feet, swells 5 ft, air temperature 13°C, max. heave 0.5 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 3 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Cores U1362A-7R and 8R arrived on deck at 0420 and 0910 hr, respectively, advancing to 399 m below seafloor (mbsf). During the cutting of Core 9R, the drill string got stuck at 1400 hr. The pipe was worked free between 1400 and 1545 hr, and the core arrived on deck at 1620 hr. The hole was washed, reamed, and swept with mud and the next core barrel was deployed at 2045 hr. Coring proceeded to 412 mbsf, at which point the pipe got stuck again. Core recovery has averaged to 23% so far.
Shipboard samples were taken from the first four cores. Core description and lab measurements are in progress. Several scientists and outreach officers participated in a videoconference with a SeaWorld (Florida) science camp. Weather on site: winds 15 knots gusting to 18 knots, seas 5 feet, swells 5 ft, air temperature 13.5°C, max. heave 1.3 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 2 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Cores U1362A-5R and 6R arrived on deck at 0230 and 0900 hr, respectively, advancing to 380 mbsf. The hole was conditioned with several mud sweeps and the next core barrel was deployed at 1345 hr. More problems were encountered, the empty core barrel was recovered, and the hole was again washed, reamed, and swept with mud between 1600 and 2045 hr. The next core barrel was deployed at 2045 hr and the cutting of Core 7R proceeded until midnight to a depth of 382 mbsf.
Cores 5R-6R were curated and measured with the whole-round loggers. The petrologists started describing Core 2R and a meeting was held at noon to discuss shipboard samples based on this preliminary examination. The decision was made to defer shipboard sampling by a day. Outreach officers continued working on several projects. Weather on site: winds 22 knots gusting to 26 knots, seas 7 feet, swells 6 ft, air temperature 14°C, max. heave 1.5 m, max. pitch 1°, max. roll 1°.


Daily Science Report for 1 August 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Coring at Hole U1362A continued to 352.6 m below seafloor (mbsf). Core 2R arrived on deck at 0535 hr with an average rate of penetration of 2 m/hr. Core 3R was cored to 362.2 mbsf and arrived on deck at 1125 hr. Core 4R was recovered at 1445 hr after advancing only 2.5 m because the pipe got stuck. The core barrel for Core 5R was deployed at 2045 hr and coring to 370.2 continued until midnight.
Cores 2R-4R were curated and measured with the whole-round loggers. Outreach officers continued working on several projects. Weather on site: winds 20 knots gusting to 25 knots, seas 7 feet, swells 5 ft, air temperature 14°C, max. heave 1.0 m, max. pitch 2°, max. roll 2°.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Tuesday, 02-Nov-2010 15:57:52 CDT.